# Digital currency hardware wallets and the essence of money

### 28 Feb 2026 v4

@@ 1 Introduction
Some proponents of retail CBDC, including central banks, have argued that it should have cash-like properties to deliver value to end-users [ECB 2021, Bank of England 2023]. This argument is at least partially right. For example, it might be expected that a cash-like system would afford users privacy by design, resistance to discrimination and profiling, and the ability to possess and control the monetary assets that they use to participate in the economy and satisfy their economic needs. These are certainly desirable properties. However, some authorities have interpreted the argument that a digital payment system must be cash-like to mean that a digital payment system must enable “dual-offline” transactions, namely those without connectivity of any kind between either of the transacting parties and any third parties, in contrast to “single-offline” transactions, wherein a mutually trusted third party is reachable by one of the counterparties. The “dual-offline” argument constitutes a leap of logic that must be examined carefully, starting with its implicit assumptions that any new payment system must be appropriate for all scenarios, that existing solutions such as cash are insufficient to handle payment scenarios in which neither party has access to a network connection, and that there is a specific need for future digital currency infrastructure to enable a kind of digital transaction that is not commonly available to users today.
@@
However, there is another, competing view that maintains that the purpose of digital currency is to provide a digital form of cash (“e-cash”), and this view is supported by past digital currency design proposals over the past four decades. Some of the most groundbreaking systems, including not only Bitcoin [Nakamoto 2009] but also DigiCash [Pitta 1999] two decades earlier, drew attention by introducing a way to hold money outside of custodial accounts. Both of these systems were sometimes described as “cryptocurrency” because they were engineered using cryptographic techniques. Importantly, the tokens that users held were explicitly designed to be cash-like, as can be inferred from their names, and intended to be used for payments and not linked to accounts. In the case of Bitcoin, tokens were created and spent using an immutable ledger (in the form of a blockchain system), and in the case of DigiCash, tokens were created and spent as bearer instruments using a redemption process with a central operator, which is assumed to be responsible and accountable for its issuances, which can be held anonymously by prospective payers. The idea that central banks would never support such an architecture ignores the fact that central banks have always supported the issuance and circulation of tokens outside accounts, in the form of cash. Within this framework, offering cash in digital form is the next logical step. We suggest that the “e-cash” vision of digital currency is largely absent from the discussion of the design of digital currency hardware wallets, and we argue that this neglect is a mistake. Contemporary examples of “e-cash” systems with significant deployments include GNU Taler [GNUnet and Taler] and (de facto) Privacy Pass [Davidson 2018], both of which feature the use of non-custodial wallets to hold tokens for redemption.
@@
E-cash systems are designed to work without tracking the ownership of individual tokens on a ledger. A generic asset “template” containing information such as a public key is first created by a prospective consumer, who then uses a zero-knowledge protocol to request a privacy-preserving certification of the digital asset, such as a blind signature, from the issuer, which then becomes part of the asset. The issuer can now recognise the certified digital asset as valid without recognising its owner or the circumstances of its creation. The asset is stored by the end-user directly and is not recorded on a ledger. Double-spending can be prevented in a few ways, one of which is to assume (a) that the issuer processes all transactions as redemptions, (b) that the issuer maintains a ledger to keep track of the set of redeemed assets that it receives in these transactions, and (c) that the issuer, upon receiving an asset in a transaction, checks the ledger to verify that the asset has not already been redeemed [Chaum 1982, Chaum 2021]. An alternative approach, which we call oblivious management, does not require the issuer to process transactions or maintain a ledger of redeemed tokens. In such a system, the assets carry their own provenance information, linking transaction information they carry to rolled-up aggregation entries, one per time period, on a ledger [Goodell 2023, Friolo 2025]; however, neither the tokens nor the value they represent are represented on the ledger at all. In both cases, issuers might keep accounting records of the tokens that they issue or redeem, but the privacy-preserving certification unlinks the issuance of an asset, where the party requesting a new asset may be identified, from its subsequent payment transaction and redemption, wherein only the payee is known.
@@
The aforementioned considerations raise the question of how end-users will hold digital currency assets, and in particular, what the role of a wallet should be. ISO 22739:2024 defines a wallet as an “application or mechanism used to generate, manage, store or use private keys and public keys or other digital assets” [ISO 22739 2024a]. The simplicity of this definition would seem to understate the potentially complex features required to implement token management and custody, an exploration of which is given by Barresi and Zatti [Barresi 2023]. We address some of the assumptions that Barresi and Zatti make about the design of wallets and suggest different ways of viewing the problem. In particular, we critically address the arguments that dual-offline payments constitute the most important use case for digital currency (see Section 2), that trusted computing is an indispensable enabler of digital currency solutions (see Section 3), that digital currency must be held in accounts (see Section 4), and that a wallet is properly seen as “not a place where you can store cash [but] that helps you keep track of it” [Barresi 2023] (see Section 6). Additionally, we address the argument implicit to some e-cash solutions that issuance and transaction processing must be performed by the same actor (see Section 5). Throughout this article, we shall reference a set of specific designs, including e-cash designs, such as GNU Taler, that are based on bearer instruments, as well as a proposal for how to integrate bearer instruments into a mechanism for using an integrity provider to transfer digital assets obliviously, without maintaining records of the assets, without immediately redeeming them for the value they represent, and without recording any information about the assets or the counterparties involved in transactions [Goodell 2023, Friolo 2025].
@@ 2 The sophistry of dual-offline payments
In addition to the ECB [ECB 2025], several prominent organisations, including the International Monetary Fund [Kiff 2022] as well as other central banks, such as the Bank of England [Bank of England 2022] and the Bank of Canada [2], have released reports suggesting that dual-offline payments are desirable or even essential to the functionality of a future retail CBDC system. Such bold statements are striking, considering the prevailing lack of consensus about what the requirements for digital currency should be in general. In this section, we shall critically explore the argument (or implicit assumption) that the purpose of digital currency is to facilitate access to custodial accounts in a dual-offline payment scenario.
@@ 2.1 Kinds of payments
With this in mind, it is reasonable to ask: What problem do proponents of retail CBDC intend to solve, for which dual-offline payments constitute a better solution than online payments involving assets that are held offline? We acknowledge that there are some limited real-world use cases for dual-offline payments, such as payments aboard transportation systems that lack reliable network connectivity throughout the system [1]. However, such examples are rare indeed, owing to the prevalence of network connectivity in locations suitable for digital transactions. Furthermore, we note that payment scenarios similar to this example do not actually require dual-offline payments; their requirements could also be satisfied by time-shifting, in which a consumer purchases e-money tokens online and consummates transactions with a predetermined merchant offline, by selectively sending to the merchant tokens for which the consumer had already transferred control to the merchant [Goodell 2023]. (The consumer and the merchant can use a fair exchange protocol to recover any unspent tokens later.) And so, it is fair to ask the question implicit to the title of a report by the UK House of Lords, “Central bank digital currencies: a solution in search of a problem?” [House of Lords 2021].
@@ 2.2 Motivations for digital currency
We might reasonably conclude that non-custodial wallets, which allow users to hold digital assets directly, are a foundational benefit of digital currency. However, even if we accept that non-custodial wallets are essential to digital currency, the idea digital currency held offline must be spent offline is certainly false. Users have an interest to possess and control digital assets on devices of their choosing, rather than devices associated with attributable accounts, regardless of the context in which they spend them, for reasons of privacy and censorship resistance [Kahn 2005], which we shall address in Section 3. Holding digital assets directly takes custodians out of outbound transactions, allowing users to avoid the risk that the custodians might extract fees, block certain (or all) transactions from taking place, or build profiles of the behaviour of their customers by monitoring the details of their transactions, such as time, location, and counterparty. Bearer wallets can be used in a variety of contexts, including but not limited to in-person POS transactions. Importantly, consumers can realise the privacy and anti-censorship benefits of non-custodial wallets in a fully online context such as e-commerce, for example, if they are using a system that supports asymmetric privacy [Tinn 2024].
Proponents of CBDC are right to acknowledge the secular decline in the use of cash in retail consumer transactions as a primary motivator for digital currency. For example, in the UK, banks are closing ATMs and branches [Access to Cash Review 2019], making cash more difficult for consumers to withdraw and for businesses to manage cash [Tisher 2020]. Although court judgments in some jurisdictions suggest that the legal tender status of banknotes implies mandatory acceptance in all but exceptional circumstances [European Union 2021], jurisprudence in other jurisdictions only recognises legal tender as something that must be accepted in the settlement of a debt [Bank of England 2025], admitting the possibility that vendors can freely choose what forms of payment they will accept in the general case.
@@ 2.3 Hold offline, transact online
It is worth asking whether governments or central banks are harbouring an intention to do away with cash entirely, or whether the presumption of such an intention is shaping the view of what CBDC should be. The argument that a larger share of domestic transactions should be brought inside the scope of the formal economy is certainly not new [FATF 1996, Camera 2001]. Recently, the governments of certain countries, including India [Ministry of Finance 2025] and Indonesia [Septiningrum 2024], have taken steps toward eliminating cash entirely, an objective that is consistent with prevailing business-led approaches to phase out cash transactions and encourage digital transactions in general. However, achieving that goal, whether it is legitimate or not, does not require CBDC; some countries have already implemented regulations to prevent the use of cash for large transactions or to control bank deposits comprising large amounts of cash [Financial Conduct Authority 2023]. Moreover, few monetary authorities or government regulators have called for the abolition of cash, whereas many, including those of Switzerland [Reuters 2024], the United Kingdom [Bank of England 2024], and the United States [Federal Reserve 2022], have pledged to preserve access to cash for the foreseeable future. Notably, Sweden has rolled back its drive toward creating a cashless society to support cash infrastructure in a near-cashless economy, at great cost [Library of Congress]. So the idea that CBDC is mostly a justification for phasing out cash, as has been speculated by some, is not corroborated by the facts, not yet at least.
@@ 3 Certified hardware as a de facto custodian
Barresi and Zatti explicitly assume the use of trusted hardware to draw transactions into the security envelope of a trusted custodian [Barresi 2023], implicitly drawing the hardware manufacturer into the role of trusted third party. In their model, the wallet would be a means of transacting with money held in custodial accounts that would be reconciled ex post.
@@ 3.1 Fair exchange
If we define a fair exchange as a transaction mechanism that ensures that “at the end of an exchange [between two players], either each player receives the item it expects or neither player receives any additional information about the other’s item” [Asokan 1998], then it is not possible to have a fair exchange between two parties without the involvement of a third party; this is a foundational truth [Pagnia 1999]. With the exchange of (unregistered) physical objects, possession of the object itself can be sufficient to demonstrate the veracity of the exchange that has taken place. But with digital objects the situation is more complicated. It is possible to have fair exchange without connection to the Internet, provided that a mutually trusted third party is accessible locally, perhaps via a local network, as might be possible with ad hoc networks. In particular, it is not necessary for transacting parties to have connectivity to the issuer of the assets being transacted, since the mutually trusted third party might not be the issuer, and it is not necessary for transacting parties to have connectivity to providers of attributable accounts, such as banks, since the mutually trusted third party might not be an account provider or custodian. Transacting parties can provide assurance, for example, in the form of cryptographic proofs or signatures, that any applicable redemption and compliance obligations will be fulfilled.
@@ 3.2 The perils of trusted hardware
    Trusted computing introduces significant security risks particular to digital currency. The risks of relying upon the security of a hardware manufacturer are too great for a token-based CBDC context. In particular, a system design that relies upon trusted hardware components is fragile by design. If the trusted hardware fails to function as advertised or is compromised by an attacker, then users could potentially transact the same tokens an arbitrary number of times without bound. Whereas with account-based systems, such as credit and debit cards, the cost associated with the compromise of a hardware device is limited to the size of the account with which it is associated, for a privacy-respecting system in which money is not linked to identity of the payer, the system-level costs of such an attack would be huge, and the incentives would be tremendously enticing for attackers. We note that modern devices, such as debit cards, that rely upon tamper-proof hardware for the exchange of value generally interact with accounts [ISO 8583 2023]. If their security is compromised, then the value available to attackers is limited by rules and operational procedures implemented by custodians. Without custodial accounts, such backup mechanisms are not possible, and so, the consequences of a successful attack on the hardware root of trust, such as one that impacts the cryptographic algorithm implemented in hardware or the mechanism that protects it from tampering, might not be addressable by a software upgrade, in turn requiring all users to replace their devices. Therefore, if both privacy and security are objectives of a digital currency system, it is preferable to rely upon a protocol, rather than the impermeability of end-user devices, to enforce the rules necessary to achieve the required level of security. Such a protocol can be used whenever one of the parties has access to the network, and under such circumstances, it would be better to allow the use of such a protocol rather than assume that trusted hardware modules would not be breached and would guarantee that all devices operate within the security envelope of a trusted custodian. To assume that all hardware wallets must take the latter approach because some circumstances might require it is tantamount to overspecification at the expense of users who would readily hold assets offline but transact them online. It is worth noting that when Chaum, Grothoff, and Moser argue that “any economically producible device that stores tokens with monetary value in an individual’s possession, and enables offline transactions… will be the target of successful forgery attacks…” [Chaum 2021], they are referring to wallets that perform dual-offline transactions, in contrast to wallets that hold tokens offline and transact them online. While overspecification of device requirements might certainly appeal to device manufacturers seeking compulsory mandates for their proprietary technology, it is important to consider that protocols that do not rely upon a hardware root of trust can function perfectly well outside the dual-offline context. Such protocols do not depend upon the imperviousness of end-user devices to attacks by those who possess them, and such attacks tend to fall within the capabilities of adversarial state actors and global crime syndicates.
@@
    Secure elements are unaccountable to their users. Because trusted hardware is impenetrable to users, users cannot know with certainty what their devices are doing. The inability to analyse the operations of a device undermines the trust that users can and should place in the device. In essence, the device serves a second master that is not the user. Users cannot trust what they cannot verify, so they have reason to be suspicious of unchecked promises on the part of hardware manufacturers or other privileged guarantors [Stallman 2002]. For example, tamper-proof components might store and furnish information that can be used to identify the user or the user’s previous transactions. By contrast, with free hardware design111“design that permits users to use the design (i.e., fabricate hardware from it) and to copy and redistribute it, with or without changes” [Stallman 2015]., skilled members of the community, unbound by nondisclosure agreements, can conduct audits on the design of the hardware and provide reports to the public. Of course, users with the means to do so can personally conduct the audits themselves or hire contractors to conduct the audits for them. However, it is not necessary for every user to actually run an audit; evidence suggests that similar benefits are achieved so long as all users can perform audits if they choose to do so [Heydt-Benjamin 2006].

### 12 Mar 2026 v5

@@ 1 Introduction
Some proponents of retail CBDC, including central banks, have argued that it should have cash-like properties to deliver value to end-users [ECB 2021, Bank of England and HM Treasury 2023]. This argument is at least partially right. For example, it might be expected that a cash-like system would afford users privacy by design, resistance to discrimination and profiling, and the ability to possess and control the monetary assets that they use to participate in the economy and satisfy their economic needs. These are certainly desirable properties. However, some authorities have interpreted the argument that a digital payment system must be cash-like to mean that a digital payment system must enable “dual-offline” transactions, namely those without connectivity of any kind between either of the transacting parties and any third parties, in contrast to “single-offline” transactions, wherein a mutually trusted third party is reachable by one of the counterparties. The “dual-offline” argument constitutes a leap of logic that must be examined carefully, starting with its implicit assumptions that any new payment system must be appropriate for all scenarios, that existing solutions such as cash are insufficient to handle payment scenarios in which neither party has access to a network connection, and that there is a specific need for future digital currency infrastructure to enable a kind of digital transaction that is not commonly available to users today.
@@
However, there is another, competing view that maintains that the purpose of digital currency is to provide a digital form of cash (“e-cash”), and this view is supported by past digital currency design proposals over the past four decades. Some of the most groundbreaking systems, including not only Bitcoin [Nakamoto 2009] but also DigiCash [Pitta 1999] two decades earlier, drew attention by introducing a way to hold money outside of custodial accounts. Both of these systems were sometimes described as “cryptocurrency” because they were engineered using cryptographic techniques. Importantly, the tokens that users held were explicitly designed to be cash-like, as can be inferred from their names, and intended to be used for payments and not linked to accounts. In the case of Bitcoin, tokens were created and spent using an immutable ledger (in the form of a blockchain system), and in the case of DigiCash, tokens were created and spent as bearer instruments using a redemption process with a central operator, which is assumed to be responsible and accountable for its issuances, which can be held anonymously by prospective payers. The idea that central banks would never support such an architecture ignores the fact that central banks have always supported the issuance and circulation of tokens outside accounts, in the form of cash. Within this framework, offering cash in digital form is the next logical step. We suggest that the “e-cash” vision of digital currency is largely absent from the discussion of the design of digital currency hardware wallets, and we argue that this neglect is a mistake. Contemporary examples of “e-cash” systems with significant deployments include GNU Taler [GNUnet and Taler] and (de facto) Privacy Pass [Davidson et al. 2018], both of which feature the use of non-custodial wallets to hold tokens for redemption.
@@
E-cash systems are designed to work without tracking the ownership of individual tokens on a ledger. A generic asset “template” containing information such as a public key is first created by a prospective consumer, who then uses a zero-knowledge protocol to request a privacy-preserving certification of the digital asset, such as a blind signature, from the issuer, which then becomes part of the asset. The issuer can now recognise the certified digital asset as valid without recognising its owner or the circumstances of its creation. The asset is stored by the end-user directly and is not recorded on a ledger. Double-spending can be prevented in a few ways, one of which is to assume (a) that the issuer processes all transactions as redemptions, (b) that the issuer maintains a ledger to keep track of the set of redeemed assets that it receives in these transactions, and (c) that the issuer, upon receiving an asset in a transaction, checks the ledger to verify that the asset has not already been redeemed [Chaum 1983, Chaum et al. 2021]. An alternative approach, which we call oblivious management, does not require the issuer to process transactions or maintain a ledger of redeemed tokens. In such a system, the assets carry their own provenance information, linking transaction information they carry to rolled-up aggregation entries, one per time period, on a ledger [Goodell et al. 2023, Friolo et al. 2025]; however, neither the tokens nor the value they represent are represented on the ledger at all. In both cases, issuers might keep accounting records of the tokens that they issue or redeem, but the privacy-preserving certification unlinks the issuance of an asset, where the party requesting a new asset may be identified, from its subsequent payment transaction and redemption, wherein only the payee is known.
@@
The aforementioned considerations raise the question of how end-users will hold digital currency assets, and in particular, what the role of a wallet should be. ISO 22739:2024 defines a wallet as an “application or mechanism used to generate, manage, store or use private keys and public keys or other digital assets” [ISO 22739 2024a]. The simplicity of this definition would seem to understate the potentially complex features required to implement token management and custody, an exploration of which is given by Barresi and Zatti [Barresi and Zatti 2023]. We address some of the assumptions that Barresi and Zatti make about the design of wallets and suggest different ways of viewing the problem. In particular, we critically address the arguments that dual-offline payments constitute the most important use case for digital currency (see Section 2), that trusted computing is an indispensable enabler of digital currency solutions (see Section 3), that digital currency must be held in accounts (see Section 4), and that a wallet is properly seen as “not a place where you can store cash [but] that helps you keep track of it” [Barresi and Zatti 2023] (see Section 6). Additionally, we address the argument implicit to some e-cash solutions that issuance and transaction processing must be performed by the same actor (see Section 5). Throughout this article, we shall reference a set of specific designs, including e-cash designs, such as GNU Taler, that are based on bearer instruments, as well as a proposal for how to integrate bearer instruments into a mechanism for using an integrity provider to transfer digital assets obliviously, without maintaining records of the assets, without immediately redeeming them for the value they represent, and without recording any information about the assets or the counterparties involved in transactions [Goodell et al. 2023, Friolo et al. 2025].
@@ 2 The sophistry of dual-offline payments
In addition to the ECB [ECB 2025], several prominent organisations, including the International Monetary Fund [Kiff 2022] as well as other central banks, such as the Bank of England [Bank of England 2022] and the Bank of Canada [Miedema 2020], have released reports suggesting that dual-offline payments are desirable or even essential to the functionality of a future retail CBDC system. Such bold statements are striking, considering the prevailing lack of consensus about what the requirements for digital currency should be in general. In this section, we shall critically explore the argument (or implicit assumption) that the purpose of digital currency is to facilitate access to custodial accounts in a dual-offline payment scenario.
@@ 2.1 Kinds of payments
With this in mind, it is reasonable to ask: What problem do proponents of retail CBDC intend to solve, for which dual-offline payments constitute a better solution than online payments involving assets that are held offline? We acknowledge that there are some limited real-world use cases for dual-offline payments, such as payments aboard transportation systems that lack reliable network connectivity throughout the system [Clark 2018]. However, such examples are rare indeed, owing to the prevalence of network connectivity in locations suitable for digital transactions. Furthermore, we note that payment scenarios similar to this example do not actually require dual-offline payments; their requirements could also be satisfied by time-shifting, in which a consumer purchases e-money tokens online and consummates transactions with a predetermined merchant offline, by selectively sending to the merchant tokens for which the consumer had already transferred control to the merchant [Goodell et al. 2023]. (The consumer and the merchant can use a fair exchange protocol to recover any unspent tokens later.) And so, it is fair to ask the question implicit to the title of a report by the UK House of Lords, “Central bank digital currencies: a solution in search of a problem?” [House of Lords 2021].
@@ 2.2 Motivations for digital currency
We might reasonably conclude that non-custodial wallets, which allow users to hold digital assets directly, are a foundational benefit of digital currency. However, even if we accept that non-custodial wallets are essential to digital currency, the idea digital currency held offline must be spent offline is certainly false. Users have an interest to possess and control digital assets on devices of their choosing, rather than devices associated with attributable accounts, regardless of the context in which they spend them, for reasons of privacy and censorship resistance [Kahn et al. 2005], which we shall address in Section 3. Holding digital assets directly takes custodians out of outbound transactions, allowing users to avoid the risk that the custodians might extract fees, block certain (or all) transactions from taking place, or build profiles of the behaviour of their customers by monitoring the details of their transactions, such as time, location, and counterparty. Bearer wallets can be used in a variety of contexts, including but not limited to in-person POS transactions. Importantly, consumers can realise the privacy and anti-censorship benefits of non-custodial wallets in a fully online context such as e-commerce, for example, if they are using a system that supports asymmetric privacy [Tinn 2024].
Proponents of CBDC are right to acknowledge the secular decline in the use of cash in retail consumer transactions as a primary motivator for digital currency. For example, in the UK, banks are closing ATMs and branches [Access to Cash Review 2019], making cash more difficult for consumers to withdraw and for businesses to manage cash [Tisher et al. 2020]. Although court judgments in some jurisdictions suggest that the legal tender status of banknotes implies mandatory acceptance in all but exceptional circumstances [European Union 2021], jurisprudence in other jurisdictions only recognises legal tender as something that must be accepted in the settlement of a debt [Bank of England 2025], admitting the possibility that vendors can freely choose what forms of payment they will accept in the general case.
@@ 2.3 Hold offline, transact online
It is worth asking whether governments or central banks are harbouring an intention to do away with cash entirely, or whether the presumption of such an intention is shaping the view of what CBDC should be. The argument that a larger share of domestic transactions should be brought inside the scope of the formal economy is certainly not new [FATF 1996, Camera 2001]. Recently, the governments of certain countries, including India [Ministry of Finance 2025] and Indonesia [Septiningrum et al. 2024], have taken steps toward eliminating cash entirely, an objective that is consistent with prevailing business-led approaches to phase out cash transactions and encourage digital transactions in general. However, achieving that goal, whether it is legitimate or not, does not require CBDC; some countries have already implemented regulations to prevent the use of cash for large transactions or to control bank deposits comprising large amounts of cash [Financial Conduct Authority 2023]. Moreover, few monetary authorities or government regulators have called for the abolition of cash, whereas many, including those of Switzerland [Reuters 2024], the United Kingdom [Bank of England 2024], and the United States [Federal Reserve 2022], have pledged to preserve access to cash for the foreseeable future. Notably, Sweden has rolled back its drive toward creating a cashless society to support cash infrastructure in a near-cashless economy, at great cost [Library of Congress 2021]. So the idea that CBDC is mostly a justification for phasing out cash, as has been speculated by some, is not corroborated by the facts, not yet at least.
@@ 3 Certified hardware as a de facto custodian
Barresi and Zatti explicitly assume the use of trusted hardware to draw transactions into the security envelope of a trusted custodian [Barresi and Zatti 2023], implicitly drawing the hardware manufacturer into the role of trusted third party. In their model, the wallet would be a means of transacting with money held in custodial accounts that would be reconciled ex post.
@@ 3.1 Fair exchange
If we define a fair exchange as a transaction mechanism that ensures that “at the end of an exchange [between two players], either each player receives the item it expects or neither player receives any additional information about the other’s item” [Asokan 1998], then it is not possible to have a fair exchange between two parties without the involvement of a third party; this is a foundational truth [Pagnia and Gärtner 1999]. With the exchange of (unregistered) physical objects, possession of the object itself can be sufficient to demonstrate the veracity of the exchange that has taken place. But with digital objects the situation is more complicated. It is possible to have fair exchange without connection to the Internet, provided that a mutually trusted third party is accessible locally, perhaps via a local network, as might be possible with ad hoc networks. In particular, it is not necessary for transacting parties to have connectivity to the issuer of the assets being transacted, since the mutually trusted third party might not be the issuer, and it is not necessary for transacting parties to have connectivity to providers of attributable accounts, such as banks, since the mutually trusted third party might not be an account provider or custodian. Transacting parties can provide assurance, for example, in the form of cryptographic proofs or signatures, that any applicable redemption and compliance obligations will be fulfilled.
@@ 3.2 The perils of trusted hardware
    Trusted computing introduces significant security risks particular to digital currency. The risks of relying upon the security of a hardware manufacturer are too great for a token-based CBDC context. In particular, a system design that relies upon trusted hardware components is fragile by design. If the trusted hardware fails to function as advertised or is compromised by an attacker, then users could potentially transact the same tokens an arbitrary number of times without bound. Whereas with account-based systems, such as credit and debit cards, the cost associated with the compromise of a hardware device is limited to the size of the account with which it is associated, for a privacy-respecting system in which money is not linked to identity of the payer, the system-level costs of such an attack would be huge, and the incentives would be tremendously enticing for attackers. We note that modern devices, such as debit cards, that rely upon tamper-proof hardware for the exchange of value generally interact with accounts [ISO 8583 2023]. If their security is compromised, then the value available to attackers is limited by rules and operational procedures implemented by custodians. Without custodial accounts, such backup mechanisms are not possible, and so, the consequences of a successful attack on the hardware root of trust, such as one that impacts the cryptographic algorithm implemented in hardware or the mechanism that protects it from tampering, might not be addressable by a software upgrade, in turn requiring all users to replace their devices. Therefore, if both privacy and security are objectives of a digital currency system, it is preferable to rely upon a protocol, rather than the impermeability of end-user devices, to enforce the rules necessary to achieve the required level of security. Such a protocol can be used whenever one of the parties has access to the network, and under such circumstances, it would be better to allow the use of such a protocol rather than assume that trusted hardware modules would not be breached and would guarantee that all devices operate within the security envelope of a trusted custodian. To assume that all hardware wallets must take the latter approach because some circumstances might require it is tantamount to overspecification at the expense of users who would readily hold assets offline but transact them online. It is worth noting that when Chaum, Grothoff, and Moser argue that “any economically producible device that stores tokens with monetary value in an individual’s possession, and enables offline transactions… will be the target of successful forgery attacks…” [Chaum et al. 2021], they are referring to wallets that perform dual-offline transactions, in contrast to wallets that hold tokens offline and transact them online. While overspecification of device requirements might certainly appeal to device manufacturers seeking compulsory mandates for their proprietary technology, it is important to consider that protocols that do not rely upon a hardware root of trust can function perfectly well outside the dual-offline context. Such protocols do not depend upon the imperviousness of end-user devices to attacks by those who possess them, and such attacks tend to fall within the capabilities of adversarial state actors and global crime syndicates.
@@
    Secure elements are unaccountable to their users. Because trusted hardware is impenetrable to users, users cannot know with certainty what their devices are doing. The inability to analyse the operations of a device undermines the trust that users can and should place in the device. In essence, the device serves a second master that is not the user. Users cannot trust what they cannot verify, so they have reason to be suspicious of unchecked promises on the part of hardware manufacturers or other privileged guarantors [Stallman 2002]. For example, tamper-proof components might store and furnish information that can be used to identify the user or the user’s previous transactions. By contrast, with free hardware design111“design that permits users to use the design (i.e., fabricate hardware from it) and to copy and redistribute it, with or without changes” [Stallman 2015]., skilled members of the community, unbound by nondisclosure agreements, can conduct audits on the design of the hardware and provide reports to the public. Of course, users with the means to do so can personally conduct the audits themselves or hire contractors to conduct the audits for them. However, it is not necessary for every user to actually run an audit; evidence suggests that similar benefits are achieved so long as all users can perform audits if they choose to do so [Heydt-Benjamin et al. 2006].